OpenStudy (anonymous):

prove: (sin3x)(1+cot2x)=sinx How?

OpenStudy (anonymous):

Well, looking at the graph of (sin 3x)(1+cot 2x) it doesn't look much like sin x. So I'm guessing that's gonna be a hard one to prove.

OpenStudy (anonymous):

did you mean cos (2x)?

OpenStudy (anonymous):

prove (sin^3x)(1+cot^2x)=sinx

OpenStudy (anonymous):

Do you mean? \[(sin\ x)^3[1+(cot\ x)^2] = sin\ x\]

OpenStudy (anonymous):

\[(sin\ x)^3[1+(cot\ x)^2] \] \[=(sin\ x)^3 + (sin\ x)^3(cot\ x)^2\] \[= (sin\ x)[(sin\ x)^2 + \frac{(sin\ x)^2(cos\ x)^2}{(sin\ x)^2}]\] \[= (sin\ x)[(sin\ x)^2 + (cos\ x)^2]\] \[= (sin\ x)(1)\] \[= sin\ x\]